# ADHint: Adaptive Hints with Difficulty Priors for Reinforcement Learning

### v1

@@ 5.5 Further Analysis
In Table 4, we vary $w_{\text{max}}$ and observe that performance is optimal when it is set to 0.2. Both increasing and decreasing the hint ratio lead to performance degradation. As shown in Figure 5(a), a 20% hint ratio is sufficient to keep rewards within the range of 0.5–0.7, providing stable learning signals at a moderate difficulty level. Further increasing the hint ratio yields excessively high rewards for hint-rollouts, which introduces larger update bias. These results suggest that the prefix segment of reasoning trajectories is crucial for guiding the policy model toward effective reasoning.
@@ 6 Conclusion
In this work, we introduce ADHint, which explicitly incorporates difficulty into hint-based on-policy learning. Enhanced by critical components that schedule the hint ratio with respect to the sample difficulty prior, estimate relative advantages with respect to the rollout difficulty posterior, and modulate token-level gradients based on distribution consistency, ADHint surpasses existing methods of comparable scale across various benchmarks, showcasing advanced reasoning abilities and out-of-distribution generalization. Despite its impressive performance, ADHint’s scaling behavior on larger models with 32B or more parameters, as well as its generalization to tasks without verifiable answers, merits further investigation.
[…]
Supplementary Material
Our supplementary material is organized as follows:
  • Section A presents the detailed workflow of ADHint.
  • Section B provides a detailed discussion of related work.
  • Section C provides additional details on training data construction, implementation details, and the reimplementation of baselines.
  • Section D offers an in-depth comparison between ADHint and the baselines, analyzing the sources of their performance differences.
  • Section E presents several case studies that provide a more intuitive view of ADHint’s behavior.
A Training Algorithm of ADHint
@@
     Compute token-wise factors $k_{i,t}$ by Eq. 5
@@
B Detailed Related Work
@@
Group Relative Policy Optimization (GRPO) [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages through intra-group sampling. This approach has become a cornerstone for the development of recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards]. In parallel, the multimodal community has explored various strategies to enhance the reasoning ability of MLLMs. Some works refine the reasoning style by encouraging models to “think with images” [zheng2025deepeyes, su2025pixel], explicitly grounding reasoning in visual features and directing attention to specific objects or regions. VL-ReThinker [wang2025vl] introduces an explicit reflection phase into the reasoning process, allowing the model to revisit and revise its initial thoughts. Other methods inject noise into rollouts [liu2025noisyrollout, wang2025perception] to improve training robustness and mitigate hallucinations. There are also approaches that design more precise reward functions for tasks such as object detection and semantic segmentation [liu2025visionreasoner, liu2025visual], as well as methods that select or curate training data to improve efficiency [li2025truth, liang2025modomodomultidomaindatamixtures]. However, these advances essentially remain optimizations of on-policy RL, primarily unlocking the latent potential of pretrained models, and they often fail to expand the models’ fundamental capability boundary.
Reinforcement Learning with Off-Policy Data
@@
Recent work has investigated how to incorporate off-policy data into RL-based post-training. Some approaches directly integrate off-policy data into the SFT or RL objective [yan2025learning, ma2025learning, dong2025rl, fu2025srft], but they struggle with the substantial distribution mismatch between off-policy data and the current policy model. To alleviate this issue, some methods design corrections for importance sampling [yan2025learning, dong2025rl], control the trade-off between SFT and RL losses via entropy-based schedules [fu2025srft], or alternate between RL and SFT phases [ma2025learning]. Hint-based methods address this problem more flexibly by guiding the model’s reasoning with partial hints extracted from off-policy data while preserving its exploration behavior. UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ the same temporally annealed hint ratios for all samples. Other methods reserve hints only for the hardest samples and determine the hint ratio from a fixed, predefined set via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo]. StepHint [zhang2025stephint] samples rollouts under multiple hint ratios and merges the resulting hint-rollouts, naive-rollouts, and ground truth off-policy data into a single group for advantage estimation. However, these methods overlook difficulty as a key factor, leading to insufficient and unstable training. In this work, we propose ADHint, which explicitly incorporates the sample difficulty prior into the hint-ratio schedule and the rollout difficulty posterior into relative-advantage estimation. This difficulty-aware design stabilizes training, enhances exploration, and expands the model’s knowledge boundary when learning from off-policy hints.
@@
C Detailed Experimental Settings
C.1 Training data
For MLLM training, we construct an off-policy hint dataset based on ViRL39K [wang2025vl], which covers STEM, chart reasoning, document reasoning, and spatial reasoning. We first remove samples that involve multi-image reasoning, text-only reasoning, or format errors, obtaining roughly 34K examples. Figure 7(a) shows the avg@8 distribution of Qwen2.5-VL-7B on this filtered set. We then sample hints with Qwen3-VL-30B-A3B-Thinking, and Figure 7(b) reports the distribution of examples for which this model fails to produce any correct hint among eight candidates. For these hard cases, we further sample hints with Qwen3-VL-235B-A22B-Thinking, whose avg@8 distribution is shown in Figure 7(c). Among samples with at least one correct hint, we discard those with hint length exceeding 8K tokens and randomly select one of the eight candidates as the final hint, with the resulting hint-length distribution shown in Figure 7(d). Finally, we remove samples that are excessively simple for Qwen2.5-VL-7B and obtain ViRL-Hint17k, which we will open-source. For LLM training, we sample 21K examples with a maximum hint length of 4K tokens from public datasets [yan2025learning].
C.2 Benchmarks
@@
  • We-Math — A large-scale benchmark that investigates human-like mathematical reasoning in MLLMs, decomposing problems into sub-problems and introducing multi-dimensional metrics to analyze knowledge acquisition, generalization, and rote memorization behaviours. We use its testmini split.
@@
For LLM evaluation, we employ six mathematical reasoning benchmarks that align with our training domain:
@@
  • AIME25 — The 2025 counterpart of AIME24, consisting of 30 similarly challenging problems from the 2025 AIME, which serves as an independent but distribution-matched test set.
@@
C.3 Implementation Details
As a supplement to the main text, Tables 6 and 7 summarize the training configs of the GRPO and SFT baselines, respectively, while Table 8 reports the additional configs of ADHint. Note that both GRPO and ADHint use 16 rollouts to ensure a fair comparison.
C.4 Baselines
@@
For methods that do not support MLLM training or whose code is not publicly available, we carefully reproduce them within the EasyR1 framework [zheng2025easyr1]. We note that several baselines exhibit instability and crash during training. We provide a detailed analysis in Section D.
@@
Qwen2.5-VL & Qwen3-VL max prompt length 4096
@@
global batchsize 128
micro batchsize 64
@@
Qwen2.5-Math max prompt length 1024
max response length 4096
@@
global batchsize 128
micro batchsize 64
num rollouts 16
epoch 1
Table 6: Training configs for GRPO baselines.
@@
min pixels 2562828256*28*28
max pixels 128028281280*28*28
Table 7: Training configs for SFT baseline.
@@
Table 8: Additional training configs for ADHint.
C.5 Reimplementation Details of Baselines
@@
D Comparative Analysis of ADHint and Baselines
As shown in Table 5 and Figure 2, ADHint consistently outperforms all baselines. Here, we provide a detailed comparative analysis of ADHint and these baselines to better understand the sources of their performance differences.
@@ D Comparative Analysis of ADHint and Baselines
Compared with pure GRPO, the $\mathrm{SFT}\rightarrow\mathrm{RL}$ pipeline improves avg@8 by 1.3% but reduces pass@1 by 12.2% at the 7B scale. This pattern indicates that SFT injects new knowledge into the model and raises the upper bound of its performance when multiple samples are allowed, whereas the subsequent RL phase, starting from an SFT-initialized policy, tends to lose part of the exploration capability that is crucial for strong pass@1 performance compared with directly training with on-policy RL. In practice, SFT encourages the model to follow a specific response style, which makes the response length and entropy during RL training concentrate in a relatively narrow range, thereby limiting exploration across diverse reasoning trajectories. In contrast, ADHint achieves the best results on both pass@1 and avg@8. By explicitly accounting for difficulty, ADHint assigns an appropriate hint ratio to each sample and estimates fair relative advantages for each rollout, balancing exploration and imitation. As a result, the policy model simultaneously strengthens its reasoning ability and absorbs external knowledge from off-policy hints.
@@
GHPO [liu2025ghpo] and HintGRPO [Huang_2025_ICCV] share a similar design philosophy: they add hints with fixed ratios (e.g., 25%, 50%) only to the hardest problems. The key difference is that GHPO injects hints into the prompt and excludes them from the training loss, whereas HintGRPO treats hints as part of the response and includes them in the loss computation. In our experiments, we observe that GHPO exhibits large update magnitudes and highly unstable training, likely because the model is not truly learning the knowledge contained in the hints but instead adapting to an out-of-distribution reasoning pattern. In contrast, HintGRPO trains more stably and maintains a high average reward during training. However, as shown in Table 5, HintGRPO performs poorly on the benchmarks and shows limited generalization. The progressive hinting strategy for the hardest samples helps obtain positive rollouts, but it ignores sample difficulty and can provide excessive guidance, causing SFT-like overfitting to the hint distribution and weakening exploration. As a result, the model attains inflated training rewards under heavy hinting but fails to translate these gains into improvements on the benchmarks. By comparison, ADHint leverages a sample difficulty prior to assigning appropriate guidance to each example and uses low-variance update signals to gradually push the model’s capability boundary, enabling it to learn from off-policy data while preserving its inherent exploration.
@@
StepHint [zhang2025stephint] mixes hint-rollouts with multiple fixed hint ratios, naive-rollouts, and ground-truth trajectories into a single group for advantage estimation without distinguishing among them. As shown in Figure 1(b) of the main text, this imbalanced intra-group estimation allows longer hint-rollouts with more positive trajectories to receive disproportionately large advantages, which biases policy updates toward imitating hints and leads to abnormal update behavior. This issue becomes more severe on ViRL-Hint17k, where the average hint length is about 3,100 tokens, while Qwen2.5-VL models typically generate only 400–500 tokens. In contrast, ADHint introduces AE-RDP, which estimates rollout advantages using the difficulty posterior evaluated from naive-rollouts and hint-rollouts, balancing learning between the two rollout types. CGM and Selective Masking further modulate token-level gradients within hints, which promotes more stable and reliable policy updates.
E Case Study
Figure 8 shows a representative training example. This case includes the extra prompt we add beyond the user query, which instructs the model to think before producing its response. For clarity, we omit this additional prompt in the subsequent cases. We see that the naive-rollout of the policy model fails to produce the correct answer. However, when we give the policy a hint taken from the complete off-policy reasoning trajectory, it continues the reasoning process and computes the final answer $z=2\sqrt{15}$. In this way, the model not only learns the essential knowledge required to solve the problem from the hint, but also improves its reasoning ability.

### v2

@@ 5.5 Further Analysis
In Table 7, we vary $w_{\text{max}}$ and observe that performance is optimal when it is set to 0.2. Both increasing and decreasing the hint ratio lead to performance degradation. As shown in Figure 4(c), a 20% hint ratio is sufficient to keep rewards within the range of 0.5–0.7, providing stable learning signals at a moderate difficulty level. Further increasing the hint ratio yields excessively high rewards for hint-rollouts, which introduces larger update bias. These results suggest that the initial prefix of the reasoning trajectory is critical for steering the policy model toward effective reasoning.
@@ 6 Conclusion
In this work, we propose ADHint, which explicitly incorporates difficulty into hint-based on-policy learning. Enhanced by critical components that schedule the hint ratio with respect to the sample difficulty prior, estimate relative advantages with respect to the rollout difficulty posterior, and modulate token-level gradients based on distribution consistency, ADHint surpasses existing methods across diverse settings, showcasing advanced reasoning abilities and OOD generalization. Despite its impressive performance, ADHint’s scaling behavior on larger models with 32B or more parameters, as well as its generalization to tasks without verifiable answers, merits further investigation.
A Overview
Our appendix is organized as follows:
  • Section B presents the detailed workflow of ADHint.
  • Section C provides an extended discussion of related work.
  • Section D offers additional details on training data construction, experimental settings, and the reimplementation of baselines.
  • Section E conducts an in-depth comparison between ADHint and the baselines, analyzing the sources of their performance differences.
  • Section F provides several case studies to offer a more intuitive view of ADHint’s behavior.
B Training Algorithm of ADHint
@@
     Compute token-wise gradient modulation factors $k_{i,t}$ by Eq. 5
@@
C Detailed Related Work
@@
GRPO [guo2025deepseek, shao2024deepseekmath] removes the critic network from PPO [schulman2017proximal] and estimates advantages through intra-group sampling. This approach has become a cornerstone for the development of recent reasoning models [yang2025qwen3, zeng2025glm, team2025kimi, team2025longcat, ye2025limo, wang2025reinforcement, tan2025towards]. In parallel, the multimodal community has explored various strategies to enhance the reasoning ability of MLLMs. Some works refine the reasoning style by encouraging models to “think with images” [zheng2025deepeyes, su2025pixel], explicitly grounding reasoning in visual features and directing attention to specific objects or regions. VL-ReThinker [wang2025vl] introduces an explicit reflection phase into the reasoning process, allowing the model to revisit and revise its initial thoughts. Other methods inject noise into rollouts [liu2025noisyrollout, wang2025perception] to improve training robustness and mitigate hallucinations. There are also approaches that design more precise reward functions for tasks such as object detection and semantic segmentation [liu2025visionreasoner, liu2025visual], as well as methods that select or curate training data to improve efficiency [li2025truth, liang2025modomodomultidomaindatamixtures]. However, these advances primarily unlock the latent potential of pretrained models, often failing to expand their fundamental capability boundaries.
Reinforcement Learning with Off-Policy Hints
@@
Recent work has investigated how to incorporate off-policy data into RL-based post-training. Some approaches directly integrate off-policy data into the SFT or RL objective [yan2025learning, ma2025learning, dong2025rl, fu2025srft], but they struggle with the substantial distribution mismatch between off-policy data and the current policy model. To alleviate this issue, some methods design corrections for importance sampling [yan2025learning, dong2025rl], control the trade-off between SFT and RL losses via entropy-based schedules [fu2025srft], or alternate between RL and SFT phases [ma2025learning]. Hint-based methods address this problem more flexibly by guiding the model’s reasoning with partial hints extracted from off-policy data while preserving its exploration behavior. UFT [liu2025uft] and Prefix-RFT [huang2025blending] employ the same temporally annealed hint ratios for all samples. SEELE [li2025staying] introduces learnable parameters to predict instance-specific hint ratios in a multi-round training scheme. Other methods reserve hints only for the hardest samples and determine the hint ratio from a fixed, predefined set via uniform sampling [chen2025data], binary search [zhang2025bread], or progressive search [Huang_2025_ICCV, liu2025ghpo]. StepHint [zhang2025stephint] samples rollouts under multiple hint ratios and merges the resulting hint-rollouts, naive-rollouts, and ground truth off-policy data into a single group for advantage estimation. However, these methods often overlook difficulty as a key factor, leading to insufficient and unstable training. In this work, we propose ADHint, which explicitly incorporates the sample difficulty prior into the hint-ratio schedule and the rollout difficulty posterior into relative-advantage estimation. This difficulty-aware design stabilizes training, enhances exploration, and expands the model’s knowledge boundary when learning from off-policy hints.
D Detailed Experimental Settings
D.1 Training data
For MLLM training, we construct an off-policy hint dataset based on ViRL39K [wang2025vl], which covers STEM, chart reasoning, document reasoning, and spatial reasoning. We first remove samples that involve multi-image reasoning, text-only reasoning, or format errors, obtaining roughly 34K examples. Figure 5(a) shows the avg@8 distribution of Qwen2.5-VL-7B on this filtered set. We then sample hints with Qwen3-VL-30B-A3B-Thinking, and Figure 5(b) reports the distribution of examples for which this model fails to produce any correct hint among eight candidates. For these hard cases, we further sample hints with Qwen3-VL-235B-A22B-Thinking, whose avg@8 distribution is shown in Figure 5(c). Among samples with at least one correct hint, we discard those with hint length exceeding 8K tokens and randomly select one of the eight candidates as the final hint, with the resulting hint-length distribution shown in Figure 5(d). Finally, we remove samples that are excessively simple for Qwen2.5-VL-7B and obtain ViRL-Hint17k, which we will open-source.
For Medical VQA, we similarly employ Qwen3-VL-30B-A3B-Thinking to sample hints from the PMC-VQA training set [zhang2023pmc], ultimately constructing a filtered subset of 15K samples. For LLM training, we utilize NuminaMath-S curated by GHPO [liu2025ghpo], which comprises 18.3K samples combined from MATH [hendrycks2measuring] and NuminaMath-1.5 [hendrycksmath2021].
@@ D.1 Training data
D.2 Benchmarks
@@
  • We-Math — A large-scale benchmark that investigates human-like mathematical reasoning in MLLMs, decomposing problems into sub-problems and introducing multi-dimensional metrics to analyze knowledge acquisition, generalization, and rote memorization behaviours. We use the testmini split.
@@
For Medical VQA, we evaluate on the PMC-VQA test set, a comprehensive benchmark derived from PubMed Central that assesses a model’s ability to comprehend medical images and generate clinically grounded answers.
For LLM evaluation, we employ five mathematical reasoning benchmarks that align with our training domain:
@@
D.3 Implementation Details
As a supplement to the main text, Tables 9 and 10 summarize the training configs of the GRPO and SFT baselines, respectively. Table 11 reports the default configs for ADHint, with the exception of MiMo-VL-7B-SFT-2508, Qwen2.5-Math-7B, and Medical VQA (based on Qwen2.5-VL-7B), where the maximum hint ratio $w_{\text{max}}$ is set to 0.4. Note that both GRPO and ADHint use 16 rollouts to ensure a fair comparison.
D.4 Baselines
@@
For methods that do not support MLLM training or whose code is not publicly available, we carefully reproduce them within the EasyR1 framework [zheng2025easyr1]. We note that several baselines exhibit instability and crash during training. We provide a detailed analysis in Section E.
Table 9: Training configs for GRPO baselines.
@@
Qwen2.5-VL & Qwen3-VL & MiMo-VL-SFT max prompt length 4096
@@
rollout batchsize 128
global batchsize 64
@@
Qwen2.5-Math max prompt length 2048
max response length 2048
@@
lr schedule cosine
warmup ratio 0.1
rollout batchsize 128
global batchsize 32
epoch 5
Table 10: Training configs for SFT baselines.
@@
min pixels $256\times 28\times 28$
max pixels $1280\times 28\times 28$
Table 11: Training configs for ADHint.
@@
D.5 Reimplementation Details of Baselines
@@
E Comparative Analysis of ADHint and Baselines
As shown in Figure 2, ADHint consistently outperforms all baselines. Here, we provide a detailed comparative analysis of ADHint and these baselines to better understand the sources of their performance differences.
@@ E Comparative Analysis of ADHint and Baselines
As shown in Table 5.2, compared with pure GRPO, the $\mathrm{SFT}\rightarrow\mathrm{RL}$ pipeline improves avg@8 by 1.3% but reduces pass@1 by 12.2% at the 7B scale. This pattern indicates that SFT injects new knowledge into the model and raises the upper bound of its performance when multiple samples are allowed, whereas the subsequent RL phase, starting from an SFT-initialized policy, tends to lose part of the exploration capability that is crucial for strong pass@1 performance compared with directly training with on-policy RL. In practice, SFT encourages the model to follow a specific response style, which makes the response length and entropy during RL training concentrate in a relatively narrow range, thereby limiting exploration across diverse reasoning trajectories. In contrast, ADHint achieves the best results on both pass@1 and avg@8. By explicitly accounting for difficulty, ADHint assigns an appropriate hint ratio to each sample and estimates fair relative advantages for each rollout, balancing exploration and imitation. As a result, the policy model simultaneously strengthens its reasoning ability and absorbs external knowledge from off-policy hints.
@@
GHPO [liu2025ghpo] and HintGRPO [Huang_2025_ICCV] share a similar design philosophy: they add hints with fixed ratios (e.g., 25%, 50%) only to the hardest problems. The key difference is that GHPO injects hints into the prompt and excludes them from the training loss, whereas HintGRPO treats hints as part of the response and includes them in the loss computation. In our experiments, we observe that GHPO exhibits large update magnitudes and highly unstable training, likely because the model is not truly learning the knowledge contained in the hints but instead adapting to an out-of-distribution reasoning pattern. In contrast, HintGRPO trains more stably and maintains a high average reward during training. However, as shown in Table 5.2, HintGRPO performs poorly on the benchmarks and shows limited generalization. The progressive hinting strategy for the hardest samples helps obtain positive rollouts, but it ignores sample difficulty and can provide excessive guidance, causing SFT-like overfitting to the hint distribution and weakening exploration. As a result, the model attains inflated training rewards under heavy hinting but fails to translate these gains into improvements on the benchmarks. By comparison, ADHint leverages a sample difficulty prior to assigning appropriate guidance to each example and uses low-variance update signals to gradually push the model’s capability boundary, enabling it to learn from off-policy data while preserving its inherent exploration.
@@
Methods like StepHint [zhang2025stephint] mix hint-rollouts with multiple fixed hint ratios, naive-rollouts, and ground-truth trajectories into a single group for advantage estimation without distinguishing among them. As shown in Figure 1(b) of the main text, this imbalanced intra-group estimation allows longer hint-rollouts with more positive trajectories to receive disproportionately large advantages, which biases policy updates toward imitating hints and leads to abnormal update behavior. This issue becomes more severe on ViRL-Hint17k, where the average hint length is about 3,100 tokens, while Qwen2.5-VL models typically generate only 400–500 tokens. In contrast, ADHint introduces AE-RDP, which estimates rollout advantages using the difficulty posterior evaluated from naive-rollouts and hint-rollouts, balancing learning between the two rollout types. CGM and Selective Masking further modulate token-level gradients within hints, which promotes more stable and reliable policy updates.
F Case Study
Figure 7 shows a representative training example. This case includes the extra prompt we added beyond the user query, which instructs the model to think before producing its response. For clarity, we omit this additional prompt in the subsequent cases. We see that the naive-rollout of the policy model fails to produce the correct answer. However, when we give the policy a hint taken from the complete off-policy reasoning trajectory, it continues the reasoning process and computes the final answer $z=2\sqrt{15}$. In this way, the model not only learns the essential knowledge required to solve the problem from the hint but also improves its reasoning ability.